OpenStudy (anonymous):

solve 3+sqrtz-11 = sqrtz+4

jhonyy9 (jhonyy9):

us parantheses please

OpenStudy (anonymous):

ummm what?

OpenStudy (asnaseer):

is your equation like this?\[3+\sqrt{z-11}=\sqrt{z+4}\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so just square both sides to get:\[(3+\sqrt{z-11})^2=(\sqrt{z+4})^2\]\[9+6\sqrt{z-11}+z-11=z+4\]which simplifies to:\[6\sqrt{z-11}=6\]\[\sqrt{z-11}=1\]square both sides to get:\[z-11=1\]\[z=12\]